Three charges, each 5.0x10^-6 are situated at the corners of an equilateral triangle of the side 1.0m. Find the force that charges q3 experienced due to the other two charges with an angle of 30degree


1
Expert's answer
2021-03-02T18:09:59-0500
"f=k\frac{q^2}{r^2}\\f=9\cdot10^9\frac{(5\cdot10^{-6})^2}{1^2}=0.225\ N"

"F=\sqrt{2f^2-2f^2\cos{60}}=f=0.225\ N"